### v1

# Instructions for *ACL Proceedings

Abstract
Process mining aims to discover, monitor and optimize the actual behaviors of real processes. While prior work has mainly focused on extracting procedural action flows from instructional texts, rule flows embedded in business documents remain underexplored. To this end, we introduce a novel annotated Chinese dataset, BPRF, which contains 50 business process documents with 326 explicitly labeled business rules across multiple domains. Each rule is represented as a <Condition, Action> pair, and we annotate logical dependencies between rules (sequential, conditional, or parallel). We also propose ExIde, a framework for automatic business rule extraction and dependency relationship identification using large language models (LLMs). We evaluate ExIde using 12 state-of-the-art (SOTA) LLMs on the BPRF dataset, benchmarking performance on both rule extraction and dependency classification tasks of current LLMs. Our results demonstrate the effectiveness of ExIde in extracting structured business rules and analyzing their interdependencies for current SOTA LLMs, paving the way for more automated and interpretable business process automation.
Instructions for *ACL Proceedings
Anonymous ACL submission
1 Introduction
In modern enterprises, business rules codify essential policies and procedures that govern automated decision-making in domains such as finance, healthcare, and e-commerce nelson2014business; yu2013modeling; kung2007fruits. Traditional rule engines and decision-management systems require domain experts to manually specify these rules and their execution logic, a process that is well-known to be labor-intensive and error-prone sivasankari2020comparative. Consequently, organizations are increasingly seeking to automate routine processes and decisions to improve efficiency and productivity. Nevertheless, automating complex, interdependent business logic remains a challenging problem.
Despite advances in natural language processing (NLP) for procedural text, existing datasets and models friedrich2011process; epure2015automatic; quishpi2020extracting; qian2020approach; lopez2021declarative; DBLP:conf/acl/LiangLDJLW023; DBLP:conf/acl/RenZ0ZL23; du2024paged focus primarily on action sequences, where systems extract sequences of tasks or operations (e.g., "chop onions" → "heat oil" → "fry onions" in a recipe). These approaches-typified by datasets like PAGED dataset du2024paged or works on procedural graph extraction-model processes as chains of actions with temporal or logical transitions. However, business processes often hinge not on explicit sequences of actions, but on interdependent rules, which govern decision logic under varying conditions. For instance, in a banking scenario, the rule ‘If the currency type is USD, then select remittance type’ depends on a condition being met, rather than an explicit preceding action. Further decisions, such as verifying identity or collecting additional input, may occur in parallel or be conditionally triggered based on earlier rule outcomes. This distinction between action-based and rule-based modeling is crucial: action flows capture what is done in sequence, while rule flows express under what conditions actions occur. Yet, existing procedural modeling works largely overlook this rule-centric structure, leaving a gap in automated understanding of real-world business logic.
To address this gap, we construct a Chinese dataset for business processes based on rule flows dataset, BPRF, for supporting the research in this direction, where 50 documents drawn from diverse domains (banking, healthcare, e-commerce, manufacturing) yielding a total of 326 labeled rules. Each rule is represented as a <Condition, Action> pair, and we annotate logical dependencies between rules (sequential, conditional, or parallel). Moreover, we introduce ExIde, a framework for automatic business rule extraction and dependency relationships identification using LLMs. Given a business process document, ExIde applies LLM prompting strategies to extract business rules and classify the dependency type between them. We evaluate ExIde using a suite of 12 state-of-the-art LLMs, benchmarking performance on both rule extraction and dependency classification tasks for current SOTA LLMs [^1] We will release our annotated datasets after the paper acceptance..
In summary, our contributions are as follows:
  • [leftmargin=*, itemindent=0pt]
  • A novel annotated dataset of 50 business process documents containing 326 explicitly labeled business rules across multiple domains. Rules are formalized in <Condition, Action> format and each rule pair is annotated with its logical dependency (sequential, conditional, or parallel).
  • A new LLM-based extraction framework using prompt engineering to identify business rules in text and infer logical structure among them.
  • A comprehensive evaluation of 12 SOTA LLMs on rule extraction and dependency classification tasks by analyzing how prompt design and model selection affect performance.
2 Related Work
2.1 Procedural Graph Extraction
Prior work on procedural graph construction has mainly focused on modeling action sequences in process-oriented texts. For instance, maeta2015framework applied named entity recognition techniques to extract procedural concepts from texts (e.g., recipes), and constructed process graphs by identifying action-object, tool-action, and state-transition dependencies. guo2018automatic introduced a method tailored to Chinese emergency response plans, leveraging structured text formats and syntactic rules to extract response conditions, roles, and task dependencies, thereby generating BPMN models. Similarly, ren2023constructing proposed a framework to extract multiple types of dependencies, including sequential, inclusion, and constraint, from unstructured procedural documents to construct enriched procedural graphs. pal2021constructing combined pre-trained language models with graph neural networks (LM-GNN) to identify action-information flows from unstructured texts, while candido2024annotated employed BiLSTM+CRF to extract entities and their interrelations from business process descriptions for model generation.
These methods mainly target the construction of process models based on action sequences, where dependencies are inferred from temporal or logical action orders. However, many real-world business scenarios involve procedural knowledge encoded not through explicit action flows, but through interdependent business rules. To this end, we propose a novel framework for extracting structured business rules and analyzing their interdependencies. Our approach first identifies discrete business rules from unstructured texts, then models rule-to-rule dependencies, enabling the construction of rule-driven procedural graphs that better capture the underlying logic of complex business processes.
2.2 LLMs for Business Process Modeling
Recent advances in LLMs have demonstrated significant potential across a range of NLP tasks, including business process modeling. bellan2022extracting were among the first to explore LLMs usage for extracting business process elements, though their approach was primarily limited to identifying sequential dependencies and lacked mechanisms to capture complex structures such as parallelism and conditional branching. du2024paged introduced a large-scale dataset for process reasoning and fine-tuned LLMs within commercial domains (e.g., restaurant services), yet their findings remain narrowly scoped, with limited generalizability to domains such as healthcare and manufacturing.
To overcome these limitations, we present an LLMs-based method for structured business rule extraction and dependency modeling. Unlike prior work focused on linear process flows, our approach effectively identifies and represents parallel and conditional relations among rules, addressing the challenges of non-linear process structures. Additionally, we coustruct a cross-domain dataset, encompassing e-commerce, healthcare, manufacturing scenarios, etc.
3 BPRF Dataset
In this section, we first provide the related definitions in business texts and then detail the components of our dataset. Next, we outline the dataset construction methodology and conduct quality assessments to ensure reliability and consistency.
[…]
currency types including RMB, USD, JPY, GBP, and HKD. choose the corresponding cash or remittance type
The cash or remittance type includes cash and remittance the customer needs to provide the purchase amount
Table 1: An example of a business text annotation, where the Condition is described using <Slot Type, Logical Jedgement, Reference Value>.
3.1 Preliminary
Definition 1 A business text is a natural language description of a business scenario or process. It is often used to help understand and standardize business operations and management requirements in a specific domain.
Definition 2 Business rules are explicit instructions describing the conditions, actions and restrictions that must be followed in various business scenarios. A business text may contain multiple business rules. We can represent these business rules as binary sets in the form of <Condition, Action>. In addition to individual rule representation, it is important to capture the dependency relationships between them, as they reflect logical or sequential connections that govern execution of a business process. For example, the execution of a business rule may depend on the completion of another.
Hence, based on the above definitions, our dataset annotation schema primarily consists of three primary components:
Definition 3 Condition specifies a trigger for a business rule and has 3 key attributes: <Slot Type, Logical Judgement, Reference Value>, where
  1. [leftmargin=*, itemindent=0pt]
  2. *
    Slot Type represents a category of user input that corresponds to specific information types such as currency type and customer type.
  3. *
    Logical Judgement defines the logical relation between the slot type and the reference value, with permissible values including contains, equal to, less than, greater than, less than or equal to and greater than or equal to.
  4. *
    Reference Value denotes the data associated with the slot type in the business text and can be classified into two types:
    1. [leftmargin=*, itemindent=0pt]
    2. (a)
      Enumeration type, representing discrete values (e.g., specific currency types such as USD, RMB, EUR).
    3. (b)
      Numeric type, used for numerical comparisons, where a reference value (e.g., \$50,000) is subject to logical judgments such as greater than or less than.
Definition 4 Action specifies business operation to be executed once the predefined Condition are met. Moreover, when the business process ends, Action value of the last business rule can be None.
Definition 5 Dependency Relationship describes how one business rule is connected to or influenced by another. These dependencies are classified into three types:
  1. [leftmargin=*, itemindent=0pt]
  2. $\bullet$
    Sequential dependency, where one rule must be executed before another;
  3. $\bullet$
    Conditional dependency, where different reference values of a business rule determine which subsequent rule will be executed;
  4. $\bullet$
    Parallel dependency, where two or more business rules need to be executed simultaneously.
Illustrative examples about these dependencies can be seen in Fig. 4, 5, 6 in Appx.A.1. These dependencies are crucial for understanding the flow and logic of business processes.
Figure 1: Annotation process of the BPRF dataset, which includes 1). Data Collection combining collections from the internet and synthetic generation using ChatGPT; 2). Annotation with three domain experts; 3). Verification with another three experts and 4). Evaluation for inter-annotator agreement using ICC and IAA.
An annotated example is presented in Table 1, which demonstrates annotation of a business text span. The business rules extracted from the given text span describe the process of purchasing foreign exchange for users. When a user selects an appropriate currency, the corresponding condition in the business rule is staisfied, triggering the execution of an associated action, namely, choosing the corresponding cash or remittance type. Subsequently, the system proceeds to the next business rule, where users must specify cash or remittance type. Once this condition is met, the corresponding action is triggered, requiring customer to provide the purchase amount. In summary, two business rules can be extracted as shown in Table 1.
Furthermore, these two rules exhibit a sequential dependency, as the execution of the second rule is contingent upon the completion of the first.
3.2 Dataset Construction
Text Sentence Business Rule Token Logical Judgement Dependency Relationship
Contains Equal to Less than Greater than Less than or equal to Greater than or equal to Sequential Conditional Parallel
50 258 326 10609 205 86 1 17 16 1 157 24 33
Table 2: Statistics of the constructed dataset.
Business text curreny types of popular including RMB USD JPY
Annotator 1 B-Slot type I-Slot type O O O B-Reference value I-Reference value I-Reference value
Annotator 2 B-Slot type O O O O B-Reference value B-Reference value B-Reference value
Annotator 3 B-Slot type I-Slot type O O O B-Reference value I-Reference value I-Reference value
Table 3: An example of NER annotation for a business text span, where the tags ’B’,’I’, and ’O’ represent ’Beginning’, ’Inside’, and ’Outside’, respectively, following the standard BIO tagging scheme.
Our dataset is derived from both real business environments, including banks, e-commerce platforms, flight booking system, and other relevant domains. Additionally, we incorporate synthetic data generated via ChatGPT DBLP:conf/nips/Ouyang0JAWMZASR22. The dataset construction process is shown in Fig 1, involving the following steps:
  1. [leftmargin=*, itemindent=0pt]
  2. 1.
    Data Collection: Business texts were gathered from diverse real-world domains (e.g., banking, e-commerce). Additionally, synthetic business texts were generated using ChatGPT 3.5 to enhance dataset diversity. The prompt used for synthetic text generation is shown in Fig 7, where {Definition 1,2,3,4,5} correspond to the definitions provided in Section 3.1.
  3. 2.
    Annotation: Three domain experts with expertise in business processes and natural language processing annotated the dataset. They identified business rules, and labeled them following the definitions outlined in Section 3.1.
  4. 3.
    Verification: To ensure the quality and consistency of the annotations, we conducted a verification process where the annotations were reviewed and validated by another three experts.
The final dataset comprises 50 business texts and 326 business rules, with each text containing an average of 7 rules. Table 2 presents a detailed statistical summary of the dataset.
3.3 Dataset Analysis
To evaluate the quality of the constructed dataset, we conducted an assessment from two perspectives: (1) the quality of the generated business texts based on human judgment, and (2) the consistency of the annotated business rules with the original business texts.
Figure 2: The human evaluation results from every perspective among three annotators.
Human Evaluation of Business Text Quality: Following prior studies miller1979humanistic; DBLP:conf/acl/DuLLL24, we engaged three annotators to assess generated business texts across five key dimensions: Readability, Accuracy, Clarity, Simplicity and Usability. Annotators rated each criterion on a scale from 1 to 5. The evaluation results across these dimensions are depicted in Fig 2. To assess reliability of annotations, we computed Intraclass Correlation Coefficient (ICC) shrout1979intraclass among annotators, with a higher ICC indicating stronger inter-annotator agreement and more reliable annotations. An ICC value of 0.75 or higher is generally considered indicative of strong agreement koo2016guideline. Finally, we calculated the average ICC across all criteria, obtaining a value of 0.904, which indicates strong agreement among annotators.
Inter-Annotator Agreement for Business Rule Annotation: Standard inter-annotator agreement (IAA) artstein2017inter measures primarily focus on classification tasks or qualitative numerical annotations, making it challenging to evaluate datasets annotated with structured business rules. To address this, we transformed our dataset into a Named Entity Recognition (NER) format, allowing for standard agreement measures. An example annotation is provided in Table 3. Three NLP experts in the phase of Verification annotated the dataset using the BIO tagging scheme. We then computed Fleiss’ Kappa artstein2017inter to measure inter-annotator agreement, obtaining a score of 0.911, which indicates almost perfect agreement (values range from 0 to 1, with scores above 0.6 considered high).
4 Methodology
Figure 3: Framework overview of ExIde, which includes two stages: 1). 5 different prompt variants are designed to guide LLMs to extract business rules and 2). a dedicated prompt is designed for dependency relationship identification among the extracted business rules from the stage 1.
In this section, we propose ExIde, a two-stage framework for automated business process extraction using LLMs, as illustrated in Fig 3. The framework includes: (1) Extraction of Business Rules and (2) Identification of Dependency Relationships. In the first stage, we design five prompt variants to guide the LLMs in extracting business rules from domain-specific texts. These prompts adopt a chain-of-thought format wei2022chain, structured according to a shared template shown in Fig 8, where Definition 1, 2, 4 refer to the definitions introduced in Section 3.1. The specific details of five prompts are as follows:
  • [leftmargin=*, itemindent=0pt]
  • Prompt 1 (Implicit Mapping): The corresponding sentences containing business rules are subsequently included in the Explain section. The corresponding business rules are then used to populate the Output section. The explanation precedes the extracted business rules, with no explicit sentence-rule alignment, thereby establishing an implicit mapping. Fig 9 provides an overview of Prompt 1 (Implicit Mapping).
  • Prompt 2 (Explicit Mapping): To enhance interpretability, this prompt explicitly links each explanatory sentence to a corresponding business rule using declarative phrasing (e.g., “This sentence corresponds to the business rule: …”). This format enforces a one-to-one mapping. The structure of Prompt 2 is illustrated in Fig 10.
  • Prompt 3 (Clarified Input): This prompt refines the input specification by explicitly naming the "Input" field in the basic template (Fig 8) as "Business Text Input". This modification aims to enhance clarity, ensuring that the input source is explicitly recognized as business-related text.
  • Prompt 4 (Logical Judgement): This prompt provides additional clarification on distinguishing business rules based on logical conditions such as ’contain’ and ’equal to’. Furthermore, it introduces a dedicated ’Note’ section to guide the handling of slot types that have multiple reference values. The details of this prompt are illustrated in Fig 11, with the corresponding content incorporated into the Logical Judgement section of the basic template.
  • Prompt 5 (Pseudo-Code): This prompt augments the extracted business rules with a pseudo-code representation to demonstrate their executable implementation. It defines functions such as select_from() for listing options and execute_action() for performing actions, thereby illustrating how business rules can be operationalized in a structured manner. Fig 12 presents the detailed design of Prompt 5, which is integrated into the Explain section of the basic template.
Prompts 3 (Clarified Input), 4 (Logical Judgement), and 5 (Pseudo-Code) are all modifications based on Prompt 1 (Implicit Mapping). These modifications were made to evaluate the models’ performance under different conditions.
In the second stage, we introduce a dedicated prompt for identifying dependency relationships among the extracted business rules. This prompt operates on the set of business rules extracted during the previous stage and is illustrated in Fig 13, where Definition 1, 2, 3, 4 refer to definitions introduced in Section 3.1. Specifically, we construct input pairs by combining extracted business rules in a pairwise manner, enabling the prompt to determine potential dependency relations between them.
5 Experiments
In this section, we present a comprehensive evaluation of LLMs for task of automatic business process extraction from our curated dataset. Specifically, we assessed 12 representative LLMs, including 5 closed-source models: GPT 4o hurst2024gpt, GPT o1 o1, ERNIE3.5 ernie3.5, QWEN2.5 qwen, and ChatGLM glm2024chatglm, where these models are based on web versions. The open-source group comprises QWEN 2.5-14b qwen2.5, Llama 3.1-8b llama, DeepSeek v3 deepseekai2024deepseekv3technicalreport, DeepSeek r1 guo2025deepseek, ERNIE-3.5-128K, GLM-4-9b-chat and QWQ-32b qwq32b. Notably, GPT o1, DeepSeek r1, and QWQ-32b are explicitly designed with advanced reasoning capabilities.
Model Prompt 1 (Implicit Mapping) Prompt 2 (Explicit Mapping) Prompt 3 (Clarified Input) Prompt 4 (Logical Judgement) Prompt 5 (Pseudo-Code) avg
Closed source
GPT 4o 0.916 0.911 0.931 0.923 0.921 0.920
GPT o1 0.903 0.894 0.932 0.899 0.915 0.909
ERNIE-3.5 0.902 0.906 0.909 0.912 0.898 0.905
QWEN2.5 0.939 0.943 0.939 0.923 0.928 0.934
ChatGLM 0.910 0.892 0.863 0.859 0.907 0.886
Open source
QWEN 2.5-14b 0.925 0.939 0.927 0.903 0.924 0.924
Llama3.1-8b 0.874 0.826 0.845 0.854 0.853 0.850
DeepSeek v3 0.934 0.927 0.943 0.93 0.905 0.928
DeepSeek r1 0.918 0.907 0.911 0.907 0.873 0.903
ERNIE-3.5-128k 0.924 0.920 0.928 0.920 0.876 0.914
GLM-4-9b-chat 0.902 0.853 0.891 0.841 0.876 0.873
QwQ-32b 0.866 0.833 0.856 0.860 0.883 0.860
avg 0.910 0.896 0.906 0.894 0.897
Table 4: NER Performances using F1 scores of 12 LLMs on 5 prompt variants, where avg indicated the average F1 scores across all prompts and all models.
Evaluation Metrics: As outlined in Section 3, task of extracting business rules is framed as a NER problem. Accordingly, we adopt F1 score as the evaluation metric to assess the performance of LLMs and prompts proposed in Sec 4, with a focus on three entity types: Slot Type, Reference Value, and Action. For the Logical Judgement component, which constitutes a multi-class classification task, we employ Macro-F1 as the evaluation metric. Similarly, the task of identifying dependency relationships between business rules is also treated as a multi-class classification problem, and is likewise evaluated using Macro-F1.
Table 4 summarizes the F1 scores of all models across five prompting strategies, each designed to elicit different aspects of model reasoning and input comprehension. We further report the average F1 score across all prompts for each model, as well as the average F1 score across all models for each prompt. The classification results for Logical Judgement and for dependency relationship identification are presented in Table 5 and Table 6, respectively. A detailed discussion of these results is provided in the following section.
NER Performance: Among all the prompt variants, Prompt 1 (Implicit Mapping) consistently achieves the highest NER performance. To illustrate this, we consider a representative example from the flight booking domain. The details about this example can be seen in Table 7.
This example exhibits a rich, hierarchical entity structure comprising categories such as flight type, travel information, and cabin class. Prompt 1 achieves the highest average NER score of 1.0 on this case, effectively guiding all LLMs to align text spans with slot types without over-constraining the logic or entity boundaries. The implicit mapping mechanism appears to mitigate ambiguity when multiple candidate spans are semantically similar (e.g., overlapping flight types or payment options), thereby enhancing span-slot alignment accuracy. Prompt 4 (Logical Judgement) and Prompt 5 (Pseudo-Code) yield lower scores, 0.9453 and 0.9483 respectively. These prompts introduce higher-order reasoning elements—Prompt 4 uses logical cues (e.g., ’contain’, ’equal to’) and notes on slot multiplicity, while Prompt 5 supplements rules with executable pseudo-code. Although such elements are potentially beneficial for rule understanding, they also increase the cognitive burden during entity recognition, potentially distracting the models from direct span-slot association and leading to reduced precision.
Furthermore, as shown in Table 4, we observe that the closed-source models generally outperform their open-source counterparts, with the exception of DeepSeek v3, which achieves a competitive average F1 score of 0.928. These results highlight the emergence of open models that rival or even surpass their closed counterparts in structured information extraction tasks.
Model Prompt 1 (Implicit Mapping) Prompt 2 (Explicit Mapping) Prompt 3 (Clarified Input) Prompt 4 (Logical Judgement) Prompt 5 (Pseudo-Code) avg
Closed source
GPT 4o 0.6379 0.5604 0.5913 0.632 0.6213 0.6086
GPT o1 0.5145 0.6315 0.6356 0.635 0.5176 0.5868
ERNIE3.5 0.3207 0.3658 0.4656 0.6379 0.3579 0.4296
QWEN2.5 0.5629 0.6465 0.5649 0.5862 0.5615 0.5844
ChatGLM 0.6114 0.4646 0.3733 0.4801 0.5494 0.4958
Open source
QWEN 2.5-14b 0.6348 0.6344 0.5708 0.5342 0.6367 0.6022
Llama3.1-8b 0.5493 0.4304 0.4089 0.5755 0.6199 0.5168
DeepSeek v3 0.6486 0.6119 0.6303 0.6457 0.6325 0.6338
DeepSeek r1 0.5565 0.5446 0.5614 0.563 0.5356 0.5522
ERNIE-3.5-128k 0.6338 0.6392 0.6386 0.6497 0.5357 0.6194
GLM-4-9b-chat 0.5912 0.5914 0.57 0.5782 0.5922 0.5846
QwQ-32b 0.6586 0.6275 0.6396 0.6192 0.6538 0.6397
avg 0.5767 0.5624 0.5542 0.5947 0.5678
Table 5: Classification results using Macro-F1 scores on Logical Judgement for 12 LLMs across 5 prompt variants, where avg indicates the average Macro-F1 scores across all prompts and all models.
Classification Performance on Logical Judgement: Among the models, QwQ-32b achieves the best average performance (0.6397), followed closely by DeepSeek v3 (0.6338) and ERNIE-3.5-128K (0.6194). These models consistently benefit from Prompt 4’s structure, suggesting stronger alignment capabilities with task-specific logic cues. Moreover, we analyze a representative case from the financial domain. The selected scenario involves a foreign currency cash deposit process at a bank, characterized by multi-step logic and conditional constraints. The text and the corresponding rules can be seen in Table 8.
For this case, the best performance was achieved by Prompt 4 (0.814), demonstrating the effectiveness of explicitly guiding the model to reason over logical expressions like “contain”,“greater than” or “equal to,” which are prevalent in this case. Its dedicated "Note" section for handling multi-valued slot types likely contributed to more precise classification. Prompt 3 achieves the second-best performance (0.546), indicating that refining input representation helps in cases with clear condition-action mappings, although the improvement is limited without dedicated logical reasoning support. The results demonstrate that logical complexity—including conditional branches based on user selections (e.g., deposit type, currency category) and regulatory constraints (e.g., amount thresholds and required documents)—poses a significant challenge for generic prompts. Prompt 4’s success stems from its dedicated reasoning design, which provides structural clarity and interpretive focus.
Closed source Open source
Model GPT 4o GPT o1 ERNIE3.5 QWEN2.5 ChatGLM QWEN 2.5 -14B Llama3.1 -8B DeepSeek v3 DeepSeek r1 ERNIE-3.5 -128K GLM-4-9B -chat QwQ-32b
Macro F1 0.6325 0.7118 0.4426 0.5481 0.3678 0.5619 0.3890 0.6385 0.7542 0.5681 0.2318 0.7161
Table 6: Classification results using Macro-F1 scores about dependency relationships of 12 LLMs.
Classification Performance on Dependency Relationships: Among closed-source models, GPT o1 achieved the highest accuracy of 0.7118, demonstrating its advanced reasoning and contextual understanding abilities. In contrast, ChatGLM lagged behind with an accuracy of 0.3678. For open-source models, QWQ-32b and DeepSeek r1 stood out with macro F1 of 0.7161 and 0.7542 respectively, surpassing many LLMs and validating their potential in reasoning-intensive tasks. These results confirm that models explicitly optimized for reasoning, such as GPT o1, DeepSeek r1, and QWQ-32b, exhibit significantly stronger performance in dependency classification.
To further examine the difference in reasoning capability, we conducted a case study on business rule dependency classification using a real-world inspired scenario: the shopping process on e-commerce platforms. The text and the corresponding rules can be seen in Table 9. The classification results obtained DeepSeek v3 are shown in Fig 14. DeepSeek v3 misclassified four out of the ten rule pairs, including sequential and conditional relationships. In contrast, DeepSeek r1 correctly identified all dependency types. This stark difference highlights the importance of reasoning ability in the identification of dependency relationships. To illustrate the superior reasoning ability of DeepSeek r1, we present its internal reasoning process seen in Fig 15. The model interprets the textual cues (e.g., ’first’, ’then’, ’next’, and ’simultaneously’) and considers whether intermediary steps exist between two rules, whether actions are meant to be concurrent, and whether dependencies imply direct execution order or not.
For instance, in Rule Pair 1 (Shopping Type → Product Information), DeepSeek v3 fails to recognize the explicitly sequential nature indicated in the text. In contrast, DeepSeek r1 correctly identifies it as a direct sequential relationship. Similarly, Rule Pairs 8 and 9 (Payment Method → Shipping Address and Payment Method → Invoice Type) are misclassified as parallel by DeepSeek v3, likely due to the word ’simultaneously’ in the text. However, DeepSeek r1 correctly infers that these are sequential dependencies—address and invoice inputs occur after the payment method is selected, even if simultaneously in practice.
This example underscores the importance of reasoning mechanisms in accurately identifying subtle and nested dependency relationships. Models like DeepSeek r1 and GPT o1, which are specifically optimized for reasoning, significantly outperform their counterparts in such scenarios.
Limitations
Appendix A Appendix
A.1 Illustrative Examples of Dependency Relationships
In this section, we provide illustrative examples of the three types of dependency relationships: sequential, conditional, and parallel. These examples aim to clarify the definitions discussed earlier, with corresponding business scenarios demonstrating the practical implications of each dependency type:
  • Sequential Dependency: Our bank supports up to 39 currency types of popular countries or regions around the world, including RMB, USD, JPY, GBP, and HKD. After selecting the appropriate currency, the customer needs to choose the corresponding cash or remittance type based on the type of business to be conducted thereafter. The cash or remittance type includes cash and remittance. Upon completing the cash/remittance type selection, the customer needs to provide the purchase amount…
  • Conditional Dependency: When choosing a car rental service, the user first needs to select the vehicle type, which can be a sedan, SUV, business vehicle, or RV. If the user chooses a sedan or SUV, they will need to further choose the rental duration, with options like 1 day, 1 week, or 1 month. If the user chooses a business vehicle or RV, they will need to choose the rental purpose, which could be for travel, road trips, or business activities. After selecting the rental duration, the user will also need to choose whether to purchase insurance…
  • Parallel Dependency: When signing up for an educational training course, the user first needs to choose the course type, such as programming, design, marketing, etc. After selecting the course type, the user needs to simultaneously choose the instruction mode and course duration. The instruction mode includes online courses and in-person courses, and the course duration can be 1 month, 3 months, or 6 months. After selecting the instruction mode and course duration, the user will need to choose the learner’s age group, which could be adults or teenagers. then, the user will proceed to the tuition payment…
Figure 4: Illustration of a sequential dependency in a banking service scenario. The process follows a strict linear order: currency type $\rightarrow$ transaction type (cash or remittance) $\rightarrow$ purchase amount. Each step is dependent on the completion of the previous one, forming a chain-like execution flow.
Figure 5: Illustration of a conditional dependency in a car rental scenario. The flow diverges based on the vehicle type selected. A sedan or SUV leads to rental duration selection, while a business vehicle or RV leads to rental purpose selection. Downstream steps, such as insurance purchase, follow the resolution of these conditional branches.
Figure 6: Illustration of a parallel dependency in an educational training scenario. After selecting the course type, the user must concurrently choose both instruction mode and course duration. These selections are independent but occur in parallel, followed by a unified downstream step to specify the learner’s age group and proceed to tuition payment.
Figures 4, 5, and 6 visually represent the business processes associated with each type of dependency.
A sequential dependency occurs when the execution of one rule depends on the completion of another rule. In other words, the first rule must be executed before the second. For example, in the case of foreign exchange services, a customer must first select a currency type (such as RMB, USD, JPY, etc.). After selecting the currency, the customer must choose the corresponding cash or remittance type (cash or remittance). Finally, after selecting the cash/remittance type, the customer is required to provide the exchange amount. Each step must occur in a prescribed order, as the actions are dependent on prior selections, as depicted in Fig. 4.
A conditional dependency is present when the next rule to be executed depends on the conditions or choices made in the previous rule. In this scenario, the execution of a rule is conditional on the reference values selected by the user. For instance, in the car rental service example, when a user selects a vehicle type, the subsequent actions depend on the choice made. If the user selects a sedan or SUV, they are prompted to choose the rental duration, with options like 1 day, 1 week, or 1 month. However, if the user selects a business vehicle or RV, they will instead need to choose the rental purpose (e.g., travel, road trip, or business activity). This dependency ensures that the process flow adjusts based on user input, as shown in Fig. 5.
A parallel dependency exists when multiple business rules are executed simultaneously. For example, when registering for an educational training course, a user first selects a course type (such as programming, design, marketing, etc.). After this selection, the user must choose both the instruction mode (online or in-person) and the course duration (1 month, 3 months, or 6 months) in parallel. Then, the learner’s age group (either adults or teenagers) is selected. Finally, the user proceeds to tuition payment after completing these selections. These choices of instruction mode and course duration occur in parallel, as illustrated in Fig. 6.
Figure 7: The prompt of synthetic text generation. This prompt enables the creation of domain-specific business texts by guiding LLMs to generate structured examples based on predefined slot types and rule patterns.
Figure 8: Template shared across the five prompt variants used in Stage 1 of ExIde. The template includes fields such as “Input,” “Explain,” and “Output,” ensuring consistency in structure while allowing each variant to introduce unique interpretative or structural features for enhanced business rule extraction.
Figure 9: The details for Prompt 1 (Implicit Mapping). This variant organizes explanations and extracted rules in separate sections without explicit alignment. The mapping between sentences and rules is implicit.
Figure 10: The details for Prompt 2 (Explicit Mapping). This variant improves interpretability by explicitly linking each explanatory sentence to its corresponding business rule through declarative statements. This enforces a one-to-one correspondence between input explanations and output rules.
Figure 11: The details for Prompt 4 (Logical Judgement). This variant enhances logical clarity by distinguishing conditions like “contain” vs. “equal to,” and adds a “Note” section to guide how to handle multiple reference values.
Figure 12: The details for Prompt 5 (Pseudo-Code). This variant supplements extracted rules with pseudo-code, demonstrating their executable form. Functions such as select_from() and execute_action() are used to translate rules into structured, programmable actions, bridging natural language and formal representation.
Figure 13: The details for the prompt of identifying dependency relationships. Given a set of extracted business rules, this prompt operates pairwise to determine whether a ’sequential’, ’conditional’, or ’parallel’ dependency exists between them. If none of these relationships exist between the rules, the output should be ’no’.
When booking a flight, passengers first need to select the flight type, including domestic flight and international flight. After confirming the flight type, passengers must fill in travel information, including departure city, destination city, and departure time. Next, passengers need to choose a cabin class, such as economy class, business class, or first class. Then, passengers must also provide the number of travelers and the payment method. The number of travelers can be single or multiple; payment methods include credit card, Alipay, and WeChat Pay.
1. < <Flight Type, includes, domestic flight, international flight>, fill in travel information> 2. < <Travel Information, includes, departure city, destination city, departure time>, choose a cabin class> 3. < <Cabin Class, includes, economy class, business class, first class>, provide the number of travelers and payment method> 4. < <Number of Passengers, includes, single, multiple>, None> 5. < <Payment Methods, includes, credit card, Alipay and WeChat Pay>, None>
Table 7: An example of flight booking domain and its corresponding business rules. The upper part provides a descriptive scenario, while the lower part translates the scenario into structured business rules in the form of <Condition, Action>.
If a customer wishes to make a deposit of foreign currency cash, they must first choose the type of foreign currency cash deposit service. Our bank mainly supports four types of services: regular current deposit, call deposit, flexible fixed-term deposit, and fixed-term lump-sum deposit. If the customer selects a regular current deposit, they must further choose the currency for the regular current deposit. The supported currencies for regular current deposits include as many as 24 foreign currencies, such as the US dollar, Japanese yen, British pound, and Hong Kong dollar. After selecting the currency, the customer needs to provide the deposit amount for the regular current deposit. If the customer selects a call deposit, flexible fixed-term deposit, or fixed-term lump-sum deposit, they must further choose the currency for these services. The supported currencies for these types include up to 9 foreign currencies, such as the Singapore dollar, US dollar, and Euro. After selecting the currency, the customer needs to provide the deposit amount for the respective service. The requirements for the deposit amount of both regular current deposits and other deposit types are as follows: If the deposit amount is less than or equal to the equivalent of USD 10,000, the customer must provide a valid identification document. If the deposit amount exceeds the equivalent of USD 10,000, the customer must provide a valid identification document along with the PRC Customs Declaration Form for Inbound Passenger Luggage with a customs stamp or a cash withdrawal receipt issued by the original deposit bank.
1. < <Business Type, includes, Regular Current Deposit, Call Deposit, Flexible Fixed-Term Deposit, Fixed-Term Lump-Sum Deposit>, None> 2. < <Business Type, equals, Regular Current Deposit>, Choose currency for regular current deposit> 3. < <Business Type, equals, Call Deposit, Flexible Fixed-Term Deposit, or Fixed-Term Lump-Sum Deposit>, Choose currency for other deposit types> 4. < <Regular Current Deposit Currency, includes, USD, JPY, GBP, HKD, up to 24 currencies>, None> 5. < <Regular Current Deposit Currency, equals, USD, JPY, GBP, HKD, etc.>, Provide deposit amount for regular current deposit> 6. < <Other Deposit Currency, includes, SGD, USD, EUR, up to 9 currencies>, None> 7. < <Other Deposit Currency, equals, SGD, USD, EUR, etc.>, Provide deposit amount for other deposit types> 8. < <Regular Current Deposit Amount, less than or equal to, equivalent of USD 10,000>, Provide valid ID document> 9. < <Regular Current Deposit Amount, greater than, equivalent of USD 10,000>, Provide valid ID document and PRC Customs Declaration Form with stamp or withdrawal receipt from original deposit bank> 10. < <Other Deposit Amount, less than or equal to, equivalent of USD 10,000>, Provide valid ID document> 11. < <Other Deposit Amount, greater than, equivalent of USD 10,000>, Provide valid ID document and PRC Customs Declaration Form with stamp or withdrawal receipt from original deposit bank>
Table 8: An example of a foreign currency cash deposit process at a bank and its corresponding business rules. The upper part provides a descriptive scenario, while the lower part translates the scenario into structured business rules in the form of <Condition, Action>.
When shopping on an e-commerce platform, the customer first needs to select a shopping type, which includes product brand, product price, and product category. After selecting the shopping type, the customer is required to provide product information, including the product name, specification, quantity, etc. Next, the customer needs to choose a payment method, which can be WeChat Pay, Alipay, or credit card. Then, the customer must simultaneously provide delivery address and invoice type. The delivery address includes province, city, street, and house number; the invoice type includes electronic invoice and paper invoice.
1. < <Shopping Type, includes, Product Brand, Product Price, Product Category>, Provide Product Information > 2. < <Product Information, includes, Product Name, Specification, Quantity, etc.>, Choose Payment Method > 3. < <Payment Method, includes, WeChat Pay, Alipay, or Credit Card>, Provide Delivery Address and Invoice Type > 4. < <Delivery Address, includes, Province, City, Street, House Number>, None > 5. < <Invoice Type, includes, Electronic Invoice and Paper Invoice>, None >
Table 9: An example of the shopping process on e-commerce platforms and its corresponding business rules. The upper part provides a descriptive scenario, while the lower part translates the scenario into structured business rules in the form of <Condition, Action>.
Figure 14: The results obtained by DeepSeek v3 in analyzing the shopping process on e-commerce platforms are presented. Pairs that were misclassified are highlighted in yellow.
Figure 15: The step-by-step reasoning process of DeepSeek r1 when analyzing the shopping process on e-commerce platforms.

### v2

@@
-Instructions for *ACL Proceedings
+Business as Rulesual: A Benchmark and Framework for Business Rule Flow Modeling with LLMs
 Abstract
@@
 Our results demonstrate the effectiveness of ExIde in extracting structured business rules and analyzing their interdependencies for current SOTA LLMs, paving the way for more automated and interpretable business process automation.
-Instructions for *ACL Proceedings
-Anonymous ACL submission
+Business as Rulesual: A Benchmark and Framework for Business Rule Flow Modeling with LLMs
+Chen Yang1, Ruping Xu1, Ruizhe Li2, Bin Cao1thanks: Corresponding author., Jing Fan1 Zhejiang University of Technology, China1 University of Aberdeen, UK 2 {yangchen,rupingxu,bincao,fanjing}@zjut.edu.cn1 ruizhe.li@abdn.ac.uk2
 1 Introduction
-In modern enterprises, business rules codify essential policies and procedures that govern automated decision-making in domains such as finance, healthcare, and e-commerce nelson2014business; yu2013modeling; kung2007fruits.
-Traditional rule engines and decision-management systems require domain experts to manually specify these rules and their execution logic, a process that is well-known to be labor-intensive and error-prone sivasankari2020comparative.
+In modern enterprises, business rules codify essential policies and procedures that govern automated decision-making in domains such as finance, healthcare, and e-commerce Nelson and Sen (2014); Yu et al.
+(2013); Küng and Hagen (2007).
+Traditional rule engines and decision-management systems require domain experts to manually specify these rules and their execution logic, a process that is well-known to be labor-intensive and error-prone Sivasankari et al.
+(2020).
 Consequently, organizations are increasingly seeking to automate routine processes and decisions to improve efficiency and productivity.
 Nevertheless, automating complex, interdependent business logic remains a challenging problem.
-Despite advances in natural language processing (NLP) for procedural text, existing datasets and models friedrich2011process; epure2015automatic; quishpi2020extracting; qian2020approach; lopez2021declarative; DBLP:conf/acl/LiangLDJLW023; DBLP:conf/acl/RenZ0ZL23; du2024paged focus primarily on action sequences, where systems extract sequences of tasks or operations (e.g., "chop onions" → "heat oil" → "fry onions" in a recipe).
-These approaches-typified by datasets like PAGED dataset du2024paged or works on procedural graph extraction-model processes as chains of actions with temporal or logical transitions.
+Despite advances in natural language processing (NLP) for procedural text, existing datasets and models Friedrich et al.
+(2011); Epure et al.
+(2015); Quishpi et al.
+(2020); Qian et al.
+(2020); López et al.
+(2021); Liang et al.
+(2023); Ren et al.
+(2023a); Du et al.
+(2024a) focus primarily on action sequences, where systems extract sequences of tasks or operations (e.g., "chop onions" → "heat oil" → "fry onions" in a recipe).
+These approaches-typified by datasets like PAGED dataset Du et al.
+(2024a) or works on procedural graph extraction-model processes as chains of actions with temporal or logical transitions.
 However, business processes often hinge not on explicit sequences of actions, but on interdependent rules, which govern decision logic under varying conditions.
@@
 In summary, our contributions are as follows:
-  • [leftmargin=*, itemindent=0pt]
   • A novel annotated dataset of 50 business process documents containing 326 explicitly labeled business rules across multiple domains.
@@
 Prior work on procedural graph construction has mainly focused on modeling action sequences in process-oriented texts.
-For instance, maeta2015framework applied named entity recognition techniques to extract procedural concepts from texts (e.g., recipes), and constructed process graphs by identifying action-object, tool-action, and state-transition dependencies.
-guo2018automatic introduced a method tailored to Chinese emergency response plans, leveraging structured text formats and syntactic rules to extract response conditions, roles, and task dependencies, thereby generating BPMN models.
-Similarly, ren2023constructing proposed a framework to extract multiple types of dependencies, including sequential, inclusion, and constraint, from unstructured procedural documents to construct enriched procedural graphs.
-pal2021constructing combined pre-trained language models with graph neural networks (LM-GNN) to identify action-information flows from unstructured texts, while candido2024annotated employed BiLSTM+CRF to extract entities and their interrelations from business process descriptions for model generation.
+For instance, Maeta et al.
+(2015) applied named entity recognition techniques to extract procedural concepts from texts (e.g., recipes), and constructed process graphs by identifying action-object, tool-action, and state-transition dependencies.
+Guo et al.
+(2018) introduced a method tailored to Chinese emergency response plans, leveraging structured text formats and syntactic rules to extract response conditions, roles, and task dependencies, thereby generating BPMN models.
+Similarly, Ren et al.
+(2023b) proposed a framework to extract multiple types of dependencies, including sequential, inclusion, and constraint, from unstructured procedural documents to construct enriched procedural graphs.
+Pal et al.
+(2021) combined pre-trained language models with graph neural networks (LM-GNN) to identify action-information flows from unstructured texts, while Candido et al.
+(2024) employed BiLSTM+CRF to extract entities and their interrelations from business process descriptions for model generation.
 These methods mainly target the construction of process models based on action sequences, where dependencies are inferred from temporal or logical action orders.
@@
 Recent advances in LLMs have demonstrated significant potential across a range of NLP tasks, including business process modeling.
-bellan2022extracting were among the first to explore LLMs usage for extracting business process elements, though their approach was primarily limited to identifying sequential dependencies and lacked mechanisms to capture complex structures such as parallelism and conditional branching.
-du2024paged introduced a large-scale dataset for process reasoning and fine-tuned LLMs within commercial domains (e.g., restaurant services), yet their findings remain narrowly scoped, with limited generalizability to domains such as healthcare and manufacturing.
+Bellan et al.
+(2022) were among the first to explore LLMs usage for extracting business process elements, though their approach was primarily limited to identifying sequential dependencies and lacked mechanisms to capture complex structures such as parallelism and conditional branching.
+Du et al.
+(2024a) introduced a large-scale dataset for process reasoning and fine-tuned LLMs within commercial domains (e.g., restaurant services), yet their findings remain narrowly scoped, with limited generalizability to domains such as healthcare and manufacturing.
 To overcome these limitations, we present an LLMs-based method for structured business rule extraction and dependency modeling.
@@
 Next, we outline the dataset construction methodology and conduct quality assessments to ensure reliability and consistency.
-[…]
-[…]
+Our bank supports up to 39 currency types of popular countries or regions around the world, including RMB, USD, JPY, GBP, and HKD.
+After selecting the appropriate currency, the customer needs to choose the corresponding cash or remittance type based on the type of business to be conducted thereafter.
+The cash or remittance type includes cash and remittance.
+Upon completing the cash/remittance type selection, the customer needs to provide the purchase amount
+Slot type Logical judgement Reference value Action
 currency types including RMB, USD, JPY, GBP, and HKD. choose the corresponding cash or remittance type
@@
 Definition 3 Condition specifies a trigger for a business rule and has 3 key attributes: <Slot Type, Logical Judgement, Reference Value>, where
-  1. [leftmargin=*, itemindent=0pt]
   1. *
@@
     Reference Value denotes the data associated with the slot type in the business text and can be classified into two types:
-    1. [leftmargin=*, itemindent=0pt]
     1. (a)
@@
 Definition 5 Dependency Relationship describes how one business rule is connected to or influenced by another. These dependencies are classified into three types:
-  1. [leftmargin=*, itemindent=0pt]
   1. $\bullet$
@@
 Our dataset is derived from both real business environments, including banks, e-commerce platforms, flight booking system, and other relevant domains.
-Additionally, we incorporate synthetic data generated via ChatGPT DBLP:conf/nips/Ouyang0JAWMZASR22.
+Additionally, we incorporate synthetic data generated via ChatGPT Ouyang et al.
+(2022).
 The dataset construction process is shown in Fig 1, involving the following steps:
-  1. [leftmargin=*, itemindent=0pt]
   1. 1.
@@
 Figure 2: The human evaluation results from every perspective among three annotators.
-Human Evaluation of Business Text Quality: Following prior studies miller1979humanistic; DBLP:conf/acl/DuLLL24, we engaged three annotators to assess generated business texts across five key dimensions: Readability, Accuracy, Clarity, Simplicity and Usability.
+Human Evaluation of Business Text Quality: Following prior studies Miller (1979); Du et al.
+(2024b), we engaged three annotators to assess generated business texts across five key dimensions: Readability, Accuracy, Clarity, Simplicity and Usability.
 Annotators rated each criterion on a scale from 1 to 5.
 The evaluation results across these dimensions are depicted in Fig 2.
-To assess reliability of annotations, we computed Intraclass Correlation Coefficient (ICC) shrout1979intraclass among annotators, with a higher ICC indicating stronger inter-annotator agreement and more reliable annotations.
-An ICC value of 0.75 or higher is generally considered indicative of strong agreement koo2016guideline.
+To assess reliability of annotations, we computed Intraclass Correlation Coefficient (ICC) Shrout and Fleiss (1979) among annotators, with a higher ICC indicating stronger inter-annotator agreement and more reliable annotations.
+An ICC value of 0.75 or higher is generally considered indicative of strong agreement Koo and Li (2016).
 Finally, we calculated the average ICC across all criteria, obtaining a value of 0.904, which indicates strong agreement among annotators.
-Inter-Annotator Agreement for Business Rule Annotation: Standard inter-annotator agreement (IAA) artstein2017inter measures primarily focus on classification tasks or qualitative numerical annotations, making it challenging to evaluate datasets annotated with structured business rules.
+Inter-Annotator Agreement for Business Rule Annotation: Standard inter-annotator agreement (IAA) Artstein (2017) measures primarily focus on classification tasks or qualitative numerical annotations, making it challenging to evaluate datasets annotated with structured business rules.
 To address this, we transformed our dataset into a Named Entity Recognition (NER) format, allowing for standard agreement measures.
@@
 Three NLP experts in the phase of Verification annotated the dataset using the BIO tagging scheme.
-We then computed Fleiss’ Kappa artstein2017inter to measure inter-annotator agreement, obtaining a score of 0.911, which indicates almost perfect agreement (values range from 0 to 1, with scores above 0.6 considered high).
+We then computed Fleiss’ Kappa Artstein (2017) to measure inter-annotator agreement, obtaining a score of 0.911, which indicates almost perfect agreement (values range from 0 to 1, with scores above 0.6 considered high).
 4 Methodology
@@
 In the first stage, we design five prompt variants to guide the LLMs in extracting business rules from domain-specific texts.
-These prompts adopt a chain-of-thought format wei2022chain, structured according to a shared template shown in Fig 8, where Definition 1, 2, 4 refer to the definitions introduced in Section 3.1.
+These prompts adopt a chain-of-thought format Wei et al.
+(2022), structured according to a shared template shown in Fig 8, where Definition 1, 2, 4 refer to the definitions introduced in Section 3.1.
 The specific details of five prompts are as follows:
-  • [leftmargin=*, itemindent=0pt]
   • Prompt 1 (Implicit Mapping): The corresponding sentences containing business rules are subsequently included in the Explain section.
@@
 In this section, we present a comprehensive evaluation of LLMs for task of automatic business process extraction from our curated dataset.
-Specifically, we assessed 12 representative LLMs, including 5 closed-source models: GPT 4o hurst2024gpt, GPT o1 o1, ERNIE3.5 ernie3.5, QWEN2.5 qwen, and ChatGLM glm2024chatglm, where these models are based on web versions.
-The open-source group comprises QWEN 2.5-14b qwen2.5, Llama 3.1-8b llama, DeepSeek v3 deepseekai2024deepseekv3technicalreport, DeepSeek r1 guo2025deepseek, ERNIE-3.5-128K, GLM-4-9b-chat and QWQ-32b qwq32b.
+Specifically, we assessed 12 representative LLMs, including 5 closed-source models: GPT 4o Hurst et al.
+(2024), GPT o1 OpenAI (2024), ERNIE3.5 Team (2023), QWEN2.5 Team (2024b), and ChatGLM GLM et al.
+(2024), where these models are based on web versions.
+The open-source group comprises QWEN 2.5-14b Team (2024a), Llama 3.1-8b AI (2024), DeepSeek v3 DeepSeek-AI and Aixin Liu (2024), DeepSeek r1 Guo et al.
+(2025), ERNIE-3.5-128K, GLM-4-9b-chat and QWQ-32b Team (2025).
 Notably, GPT o1, DeepSeek r1, and QWQ-32b are explicitly designed with advanced reasoning capabilities.
@@
 Models like DeepSeek r1 and GPT o1, which are specifically optimized for reasoning, significantly outperform their counterparts in such scenarios.
+6 Conclusion
+In this paper, we introduced a novel framework, ExIde, for extracting business rules and their interdependencies from unstructured business process documents using LLMs.
+Moreover, we constructed a comprehensive Chinese dataset, BPRF, containing 326 labeled business rules across diverse domains.
+Our experiments demonstrated the effectiveness of ExIde in accurately extracting business rules and identifying their logical dependencies using 12 state-of-the-art LLMs.
+Notably, the results highlight the importance of prompt design and reasoning-oriented model architectures in enhancing extraction accuracy.
+By bridging the gap between rule-centric business logic and LLMs, our work paves the way for more intelligent, interpretable, and automated business process modeling systems.
+Future work includes expanding the dataset, refining prompt strategies, and integrating rule execution into downstream decision support and automation systems.
 Limitations
+Despite the promising results demonstrated by ExIde and the contributions of the BPRF dataset, several limitations remain.
+First, while our dataset spans multiple domains (e.g., banking, healthcare, e-commerce), its size—326 business rules across 50 documents—may still be insufficient to capture the full diversity and complexity of business process logic in the real world.
+This could affect the generalizability of our findings.
+Second, a portion of our dataset is synthetically generated using ChatGPT, which, while enhancing coverage and diversity, may introduce distributional biases not representative of naturally occurring business texts.
+Third, our framework currently relies on prompt-based querying of LLMs, which may suffer from sensitivity to prompt formulation and inconsistencies across different model architectures and decoding parameters.
+Lastly, the current approach focuses primarily on extracting and classifying pairwise dependencies; modeling more complex global structures or constraints remains an open challenge.
+Ethical Statement
+This work presents a novel benchmark (BPRF) and framework (ExIde) for extracting business rules and their logical dependencies from business process documents.
+The dataset includes both real-world business texts collected from publicly available sources and synthetically generated texts using ChatGPT.
+All real data were obtained from publicly accessible documents and do not contain any personally identifiable information (PII) or confidential content.
+Moreover, no human subjects were involved in the model evaluation phases, and no sensitive user data were used.
+The study adheres to responsible AI development principles, including transparency, reproducibility, and minimization of potential harms.
+Our annotated dataset will be released upon paper acceptance to support further research in this area.
 Appendix A Appendix